# Digging into the ultraviolet luminosity functions of galaxies at high redshifts: galaxies evolution, reionization, and cosmological parameters

### v1

@@ Abstract
Thanks to the successful performance of the James Webb Space Telescope, our understanding of the epoch of reionization of the Universe has been advanced. The ultraviolet luminosity functions (UV LFs) of galaxies span a wide range of redshift, not only revealing the connection between galaxies and dark matter (DM) halos but also providing the information during reionization. In this work, we develop a model connecting galaxy counts and apparent magnitude based on UV LFs, which incorporates redshift-dependent star formation efficiency (SFE) and corrections for dust attenuation. By synthesizing some observations across the redshift range $4\leq z\leq 10$ from various galaxy surveys, we discern the evolving SFE with increasing redshift and DM halo mass through model fitting. Subsequent analyses indicate that the Thomson scattering optical depth was $\tau_{\rm e}=0.052^{+0.003}_{-0.002}$ and the epoch of reionization started (ended) at $z=20.58^{+6.25}_{-6.75}$ ($z=5.38^{+0.65}_{-0.70}$) which is insensitive to the choice of the truncated magnitude of the UV LFs. Incorporating additional dataset and some reasonable constraints, the amplitude of matter perturbation is found to be $\sigma_{8}=0.79\pm 0.05$, which is consistent with the standard $\Lambda$CDM model. Future galaxy surveys and the dynamical simulations of galaxy evolution will break the degeneracy between SFE and cosmological parameters, improving the accuracy and the precision of the UV LF model further.
@@ 1 Introduction
Following the formation of the first stars, the Universe emerged from its dark ages as the neutral hydrogen gas gradually became ionized under the illumination of the first lights. During the epoch of reionization, there has been debate regarding the primary energy sources responsible for rendering the universe transparent. A recent study (Jiang et al., 2022) found that the contribution of quasars was negligible, providing less than $7\%$ of the total required photons. Based on the ultra-deep James Webb Space Telescope (JWST) imaging, Atek et al. (2024) analyzed eight ultra-faint galaxies and demonstrated that the majority of the necessary photons originated from dwarf galaxies, highlighting the significant connection between reionization and galaxies. However, researches on the evolution of the neutral hydrogen fraction ($x_{\rm HI}$) across different redshifts yielded various results (Ishigaki et al., 2018; Finkelstein et al., 2019; Naidu et al., 2020), accompanied by notable uncertainties and discrepancies. For instance, Hsiao et al. (2023) imposed a constraint of $x_{\rm HI}>0.9$ at $z=10.17$ whereas Bruton et al. (2023) derived $x_{\rm HI}<0.88$ at $z=10.6$. The UV luminosity functions (UV LFs) of galaxies span a wide range of redshift from cosmic dawn to cosmic noon. Currently, the observations of UV LFs have extended up to $z\sim 15-16$ (Bouwens et al., 2023; Harikane et al., 2023; Donnan et al., 2024), offering opportunities to explore the epoch of reionization.
@@
Since galaxies were formed inside the dark matter (DM) halos, their abundance intricately tied to the distribution of DM halos (Berlind et al., 2003). The UV LFs represent the number density of galaxies at specific luminosity, serving as an observable metric that links galaxies with their DM halos. This connection unveils the underlying physical mechanisms governing galaxy formation and evolution, particularly the star formation efficiency (SFE), which quantifies the effectiveness of converting gas into stars. Based on the assumption that the most massive galaxies inhabit the most massive DM halos, the SFE of central galaxies can be derived by the abundance matching between the DM mass function and UV LFs (Bell et al., 2003; Yang et al., 2009). As DM halo masses increase, the profile of SFE reflects the influence from diversified feedback processes, including AGN feedback (Croton et al., 2006), supernovae feedback (Hopkins et al., 2012), dynamical friction effect and others (Dekel & Birnboim, 2006). Only in the local universe around $z\sim 0$, the evolution of the SFE with increasing DM halo mass has almost been ascertained (see Wechsler & Tinker (2018) for a review). However, it remains unclear whether the SFE evolves with redshift, and if it does, the nature of its evolutionary trajectory is yet to be known. Nevertheless, by establishing connections between DM halo mass and galaxy luminosity (or mass), UV LFs provide a promising avenue to extend SFE investigations to higher redshift ranges (Moster et al., 2018; Harikane et al., 2022; Wang et al., 2023).
@@
Besides exploring the epoch of reionization and the SFE, UV LFs provide additional capabilities for cosmological constraints (Yang et al., 2003; Krause & Eifler, 2017). At high redshift, UV LFs possess the ability to extract the information within Mpc scales. For instance, Sabti et al. (2022a) measured the matter power spectrum $p(k)$ at wavenumbers of $0.5\,{\rm Mpc^{-1}}<k<10\,{\rm Mpc^{-1}}$ to roughly $30\%$ precision by UV LFs at the range of $4\leq z\leq 10$. Given that the number density of DM halos is sensitive to the amplitude of mass fluctuations $\sigma_{8}$, UV LFs can be used to constrain $\sigma_{8}$ at high redshift, alleviating the $\sigma_{8}$ tension potentially (Sabti et al., 2022b). Furthermore, UV LFs provide a way to examine theoretical scenarios beyond the $\Lambda$CDM model, such as the warm DM, the ultralight axion DM and the effect of primordial non-Gaussianities (Chevallard et al., 2015; Corasaniti et al., 2017; Winch et al., 2024).
@@ 2 Method
Different from our previous work (Wang et al., 2023), which focused on calculating the DM accretion rate within the $\Lambda$CDM framework, this study employs a more comprehensive accretion model to facilitate the analysis of cosmological parameters. We utilize an alternative analytical method, the extended Press-Schechter (EPS) theory (Bond et al., 1991; Bower, 1991; Lacey & Cole, 1993), to characterize the DM accretion histories. Liu et al. (2024) conducted a comparison between two analytical models based on the EPS theory and three empirical models derived from cosmological simulations. Their findings confirmed the accuracy of the differential equation method for describing mass growth ($\dot{M}_{\rm h}={\rm d}M_{\rm EPS}/{\rm d}z$) (Correa et al., 2015). Therefore, we adopt their equations (A5-B7) for calculating mass accretion, incorporating adjustable cosmological parameters.
@@
In the redshift range of $4\leq z\leq 10$, we have collected observations of UV LFs from various surveys conducted by multiple telescopes. These include JWST (SMACS0723, GLASS, CEERS, COSMOS, NGDEEP, HUDF, CANUCS and JADES; Bouwens et al. (2023); Donnan et al. (2023); Willott et al. (2023); Adams et al. (2023b); Pérez-González et al. (2023)), the Hubble Space Telescope (HST) (the Ultra-Deep Field, the GOODS fields, the Hubble Frontier Fields parallel fields, and all five CANDELS fields; McLure et al. (2013); Finkelstein et al. (2015); Oesch et al. (2018); Bouwens et al. (2021, 2022)) , the Subaru/Hyper Suprime-Cam survey and CFHT Large Area U-band Survey (Harikane et al., 2022; Morishita & Stiavelli, 2023; Adams et al., 2023a), the Visible and Infrared Survey Telescope for Astronomy (VISTA) (COSMOS, XMM-LSS, VIDEO, and Extended Chandra Deep Field South (E-CDFS) field; Bowler et al. (2020); Adams et al. (2023a)), and the UK Infrared Telescope (UKIRT) (UKIDSS UDS field; Bowler et al. (2020)).
@@
Prior to analysis, overlaps in magnitudes, fields, and redshift bins were eliminated to prevent duplication. Nevertheless, the UV flux emitted by galaxies is invariably absorbed by interstellar dust, especially in massive galaxies and at lower redshifts. Due to this dust extinction, such observations cannot be directly integrated into the UV LFs model. Moreover, it is necessary to harmonize all observations within a consistent cosmological framework, as differing constructions of the UV LFs rely on varied cosmological parameters, specifically matter density ($\Omega_{\rm m}$) and the Hubble constant ($H_{0}$).
@@ 2.2 Eliminating the artificial biases
$\qquad\qquad\qquad+5\log_{10}\frac{(1+z)d_{{\rm L}}({\rm old})}{\rm Mpc}+25,$
@@
$\Delta M_{{\rm UV}}$ represents the bin width, and $M_{\rm UV}^{+/-}({\rm old})=M_{\rm UV}\pm\Delta M_{{\rm UV}}/2$.
@@ 2.4 Bayesian analysis
where $(x_{i},y_{i})$ are the observational data points, $f(x_{i})$ represents the predicted values, $d_{i}$ serves as a scale parameter, and $c_{i}$ is the asymmetry parameter. The priors for all primary parameters are detailed in Table 1. For an optimal balance between accuracy and efficiency, we implemented the nested sampling method, utilizing Pymultinest for Bayesian parameter estimation. We configured the analysis with 1000 live points [^2] Preliminary tests with 500 live points yielded identical results. Given the recommendation to set the number of live points higher than the dimensionality of the parameter space (Ashton et al., 2022), our configuration ensures adequate convergence. and an evidence tolerance of 0.5 to terminate the sampling process.
@@
Table 1: Prior distributions of the parameters for SFE
@@ 3.1 The Evolution of SFE and UV LFs
As shown in Figure 2, we present the best-fit UV LFs and SFE, alongside the observations. The corresponding posterior distributions are shown in Figure 8333It is worth noticing that we have tested the assumption that the SFE do not evolve with redshift ($par_{s}=0$). The (logarithmic) Bayes factor of the redshift-evolved SFE scenario compared to the non-evolving scenario is $\ln{\mathcal{B}}=473.9$, which is strongly in favor of the redshift-evolution of the SFE.. For a better view, we present the UV LFs rather than the numbers of galaxies that are used practically. It is evident that the UV LF model fits well with the observations except for the data points at the bright end, which may be attributed to the contribution from active galactic nuclear (Harikane et al., 2022) in addition to the galaxy UV LFs. The corresponding distributions of the evolved SFE are consistent with other researches (such as Fig. 2 in Wechsler & Tinker (2018)). The SFE peaks at $M_{1}\sim 10^{11.67}M_{\odot}$ for $z=4$ and $10^{13.11}\,\rm M_{\odot}$ for $z=10$ (note that the latter is only obtained for extrapolation since the extremely bright end of the UV LFs has not been detected at $z\geq 8$). At low halo masses, the profile of SFE is determined by $\beta$, which remains almost constant with $\beta\sim 0.50$ across wide redshift range. Conversely, at high masses, $\gamma$ dominates the profile which varies from $\sim 0.44$ to $\sim 1.39$, resulting in a notable steepening of the SFE. However, at $z\geq 8$, due to the absence of UV LFs at the bright end, the SFE profile at $M_{1}\geq 10^{12}\,\rm M_{\odot}$ can not be well constrained. Consequently, it remains inconclusive whether there is a reliable evolution of the SFE at high masses.
Figure 2: The left panel shows the observations of UV LFs and the best-fit results in the redshift range of $4\leq z\leq 10$. The right panel shows the corresponding distributions of SFE. The light orange region represents the range of maximum value (20%-30%) of SFE (Wechsler & Tinker, 2018). Due to the redshift-dependent nature of SFE, which connects all observations at each redshift, the estimated UV LFs are tightly constrained. Therefore, only the best-fit results are shown.
@@
where $M_{\rm trunc}$ denotes the truncation magnitude of the UV LFs. In general, $M_{\rm trunc}=-17$ corresponds to the SFR of $0.3\,\rm M_{\odot}\,yr^{-1}$ (Finkelstein et al., 2015; Bouwens et al., 2020; Oesch et al., 2018; Harikane et al., 2022). Additionally, we consider another truncated limit with $M_{\rm trunc}=-15$ following McLeod et al. (2016) and Ishigaki et al. (2018). As shown in Figure 3, the UV luminosity densities derived from the fitted UV LFs are in accordance with other observations. The empirical equation Eq.(15) in Madau & Dickinson (2014) fits well with the observations at low redshifts (i.e., $z\leq 6.5$) but tends to overestimate the UV luminosity density and the SFR density at higher redshifts.

### v2

@@ Abstract
Thanks to the successful performance of the James Webb Space Telescope, our understanding of the epoch of reionization of the Universe has been advanced. The ultraviolet luminosity functions (UV LFs) of galaxies span a wide range of redshift, not only revealing the connection between galaxies and dark matter (DM) halos but also providing the information during reionization. In this work, we develop a model connecting galaxy counts and apparent magnitude based on UV LFs, which incorporates redshift-dependent star formation efficiency (SFE) and corrections for dust attenuation. By synthesizing some observations across the redshift range $4\leq z\leq 10$ from various galaxy surveys, we discern the evolving SFE with increasing redshift and DM halo mass through model fitting. Subsequent analyses indicate that the Thomson scattering optical depth was $\tau_{\rm e}=0.054^{+0.001}_{-0.003}$ and the epoch of reionization started (ended) at $z=18.8^{+7.2}_{-6.0}$ ($z=5.3^{+0.8}_{-1.0}$) which is insensitive to the choice of the truncated magnitude of the UV LFs. Incorporating additional dataset and some reasonable constraints, the amplitude of matter perturbation is found to be $\sigma_{8}=0.80\pm 0.05$, which is consistent with the standard $\Lambda$CDM model. Future galaxy surveys and the dynamical simulations of galaxy evolution will break the degeneracy between SFE and cosmological parameters, improving the accuracy and the precision of the UV LF model further.
@@ 1 Introduction
Following the formation of the first stars, the Universe emerged from its dark ages as the neutral hydrogen gas gradually became ionized under the illumination of the first lights. Present observations have constrained the bulk of reionization to be completed by $z\sim 6$ (Fan et al., 2006), but the beginning of reionization shows less well-constraint and depends sensitively on the nature of the ionizing sources. During the epoch of reionization, there has been debate regarding the primary energy sources responsible for rendering the universe transparent. Robertson et al. (2015) analyzed the Thomson scattering optical depth $\tau$ from Planck and the abundance and luminosity distribution of distant galaxies from Hubble Space Telescope imaging. Their results strengthened the assumption that most of the ionizing photons responsible arose from high-redshift star-forming galaxies. Subsequently, Finkelstein et al. (2019) demonstrated that galaxies become more efficient producers of ionizing photons at higher redshifts and fainter magnitudes. The contribution of AGNs represented subdominant. A recent study (Jiang et al., 2022) found that the contribution of quasars was negligible, providing less than $7\%$ of the total required photons. Based on the ultra-deep James Webb Space Telescope (JWST) imaging, Atek et al. (2024) analyzed eight ultra-faint galaxies and demonstrated that the majority of the necessary photons originated from dwarf galaxies, highlighting the significant connection between reionization and galaxies. However, at $z\gtrsim 9$, the observations of JWST suggest a tension between the current data and the cosmic microwave background, since these high-redshift galaxies with $M_{\rm UV}<-15\,\rm mag$ not only can drive reionization but would end it too early (Muñoz et al., 2024). Besides, researches on the evolution of the neutral hydrogen fraction ($x_{\rm HI}$) across different redshifts yielded various results (Ishigaki et al., 2018; Finkelstein et al., 2019; Naidu et al., 2020), accompanied by notable uncertainties and discrepancies. For instance, Hsiao et al. (2024) imposed a constraint of $x_{\rm HI}>0.9$ at $z=10.17$ whereas Bruton et al. (2023) derived $x_{\rm HI}<0.88$ at $z=10.6$. The UV luminosity functions (UV LFs) of galaxies span a wide range of redshift from cosmic dawn to cosmic noon. Currently, the observations of UV LFs have extended up to $z\sim 15-16$ (Bouwens et al., 2023; Harikane et al., 2023; Donnan et al., 2024), offering opportunities to explore the epoch of reionization.
@@
Since galaxies were formed inside the dark matter (DM) halos, their abundance intricately tied to the distribution of DM halos (Berlind et al., 2003). The UV LFs represent the number density of galaxies at specific luminosity, serving as an observable metric that links galaxies with their DM halos. This connection unveils the underlying physical mechanisms governing galaxy formation and evolution, particularly the star formation efficiency (SFE), which quantifies the effectiveness of converting gas into stars. Based on the assumption that the most massive galaxies inhabit the most massive DM halos, the SFE of central galaxies can be derived by the abundance matching between the DM mass function and UV LFs (Bell et al., 2003; Yang et al., 2009). As DM halo masses increase, the profile of SFE reflects the influence from diversified feedback processes, including AGN feedback (Croton et al., 2006), supernovae feedback (Hopkins et al., 2012), dynamical friction effect and others (Dekel & Birnboim, 2006). Only in the local universe around $z\sim 0$, the evolution of the SFE with increasing DM halo mass has almost been ascertained (see Wechsler & Tinker (2018) for a review). However, it remains unclear whether the SFE evolves with redshift, and if it does, the nature of its evolutionary trajectory is yet to be unknown. Nevertheless, by establishing connections between DM halo mass and galaxy luminosity (or mass), UV LFs provide a promising avenue to extend SFE investigations to higher redshift ranges (Moster et al., 2018; Harikane et al., 2022; Wang et al., 2023).
@@
Besides exploring the epoch of reionization and the SFE, UV LFs provide additional capabilities for cosmological constraints (Yang et al., 2003; Krause & Eifler, 2017). At high redshift, UV LFs possess the ability to extract the information within Mpc scales. For instance, Sabti et al. (2022a) measured the matter power spectrum $p(k)$ at wavenumbers of $0.5\,{\rm Mpc^{-1}}<k<10\,{\rm Mpc^{-1}}$ to roughly $30\%$ precision by UV LFs at the range of $4\leq z\leq 10$. Such constraints can be derived from the cumulative stellar mass distributions (Parashari & Laha, 2023), also. Given that the number density of DM halos is sensitive to the amplitude of mass fluctuations $\sigma_{8}$, UV LFs can be used to constrain $\sigma_{8}$ at high redshift, alleviating the $\sigma_{8}$ tension potentially (Sabti et al., 2022b) (a software named GALLUMI). Furthermore, UV LFs provide a way to examine theoretical scenarios beyond the $\Lambda$CDM model, such as the warm DM, the ultralight axion DM and the effect of primordial non-Gaussianities (Chevallard et al., 2015; Corasaniti et al., 2017; Winch et al., 2024).
@@ 2 Method
Different from our previous work (Wang et al., 2023), which focused on calculating the DM mass accretion rate (MAR) within the $\Lambda$CDM framework, this study employs a more comprehensive accretion model to facilitate the analysis of cosmological parameters. We utilize an alternative analytical method, the extended Press-Schechter (EPS) theory (Bond et al., 1991; Bower, 1991; Lacey & Cole, 1993), to characterize the DM accretion histories. Liu et al. (2024) conducted a comparison between two analytical models based on the EPS theory and three empirical models derived from cosmological simulations. Their findings confirmed the accuracy of the differential equation method for describing mass growth ($\dot{M}_{\rm h}={\rm d}M_{\rm EPS}/{\rm d}z$) (Correa et al., 2015). Therefore, we adopt their equations (A5-B7) for calculating mass accretion, incorporating adjustable cosmological parameters.
@@
In the redshift range of $4\leq z\leq 10$, we have collected observations of UV LFs from various surveys conducted by multiple telescopes. These include JWST (SMACS0723, GLASS, CEERS, COSMOS, NGDEEP, HUDF, CANUCS and JADES; Bouwens et al. (2023); Donnan et al. (2023); Willott et al. (2024); Adams et al. (2024); Pérez-González et al. (2023)), the Hubble Space Telescope (HST) (the Ultra-Deep Field, the GOODS fields, the Hubble Frontier Fields parallel fields, and all five CANDELS fields; McLure et al. (2013); Finkelstein et al. (2015); Oesch et al. (2018); Bouwens et al. (2021, 2022)) , the Subaru/Hyper Suprime-Cam survey and CFHT Large Area U-band Survey (Harikane et al., 2022; Morishita & Stiavelli, 2023; Adams et al., 2023), the Visible and Infrared Survey Telescope for Astronomy (VISTA) (COSMOS, XMM-LSS, VIDEO, and Extended Chandra Deep Field South (E-CDFS) field; Bowler et al. (2020); Adams et al. (2023)), and the UK Infrared Telescope (UKIRT) (UKIDSS UDS field; Bowler et al. (2020)).
Prior to analysis, overlaps in magnitudes, fields, and redshift bins were eliminated to prevent duplication. At first, for various telescope projects, we artificially collect observation catalogues that include high redshift ($z\geq 4$) galaxies as many as possible. As listed above, most of these catalogues exhibit independence because their survey fields are completely irrelevant. Subsequently, a small fraction of the observations are removed to eliminate the effect of the non-negligible overlap. For example, at $z=9$, Pérez-González et al. (2023) derived the UV LFs from deep NIRCam observations taken in parallel with the MIRI Deep Survey of the HUDF, which has been incorporated into Bouwens et al. (2021) base on HST. In that case, the observations in Pérez-González et al. (2023) with $M_{\rm UV}<-17.92$ (the faintest observations in Bouwens et al. (2021)) are excluded. After such screening processes, all of the UV LF data we used are mutually independent.
Due to the fluctuations in the large-scale dark matter density field, cosmic variance generates the intrinsic scatter in the number density of galaxies and then introduces unavoidable errors into the UV LFs. Especially at high redshift and small survey areas, cosmic variance can be very significant. To address this, we verify whether such uncertainties have been considered in all of the UV LF observations. For the dataset that did not account for cosmic variance, we add the additional uncertainties  [^2] Given the parameter limitation of GALCV, we conservatively assume $z=5$ to calculate the cosmic variance for the observations at $z=4$. For wider survey area ($>31640\,\rm arcmin^{2}$), cosmic variance is estimated using LINCV. Both of the methods are available at https://github.com/adamtrapp/galcv. using the public Python package GALCV (Trapp & Furlanetto, 2020). For a clearer comparison across different magnitudes, redshifts, and field areas, we present the data from Bouwens et al. (2021) along with their corresponding cosmic variance in Table 4. All UV LF data used in this analysis can be accessed via GitHub  [^3] https://github.com/wangyy19/UVLF_datasets.
@@
Nevertheless, the UV flux emitted by galaxies is invariably absorbed by interstellar dust, especially in massive galaxies and at lower redshifts. Due to this dust extinction, such observations cannot be directly integrated into the UV LFs model. Moreover, it is necessary to harmonize all observations within a consistent cosmological framework, as differing constructions of the UV LFs rely on varied cosmological parameters, specifically matter density ($\Omega_{\rm m}$) and the Hubble constant ($H_{0}$).
@@ 2.2 Eliminating the artificial biases
$\qquad\qquad\qquad+5\log_{10}\frac{d_{{\rm L}}({\rm old})}{\rm Mpc}+25,$
@@
$\Delta M_{{\rm UV}}$ represents the bin width, and $M_{\rm UV}^{+/-}({\rm old})=M_{\rm UV}\pm\Delta M_{{\rm UV}}/2$. It should be noticed that the comoving volume $(V_{\rm c}\sim\frac{survey\,field}{all-sky\,filed}\times V_{\rm c}(total)$) originally represents the characteristic value that relies on the field of each galaxy surveys. Because both the model and the observations are multiplied by $V_{\rm c}$ in model fitting, the term of $\frac{survey\,field}{all\,sky\,filed}$ can be removed for logical simplification. Therefore, all of the $V_{\rm c}$ represent the value of the total-comoving volume in this work.
@@ 2.3 Redshift-dependent star formation efficiency
2.4 The effect of stochasticity
Except for the aforementioned data corrections and model-dependent assumptions, the stochastic behavior in the SFR of high-redshift galaxies fluctuates the UV LFs and increases the uncertainties of parameter estimations. Such stochasticity can be driven by the process of bursty star formation and shows an increasing tendency at higher redshift, which has been used to explain the abundance of bright galaxies at $z\gtrsim 10$ (Qin et al., 2023; Sun et al., 2023; Yung et al., 2024) that JWST detected recently. Following Mirocha & Furlanetto (2023), we add the stochasticity in accretion histories (Tacchella et al., 2018) and consider a fraction of dex log-normal scatter ($\sigma_{\rm MAR}$) in DM halo accretion rate at fixed $M_{\rm h}$. In detail, the evolution of the stochasticity with redshift increasing is assumed to follow Equation 12 with $\sigma_{{\rm MAR}_{s}}>0$, in a manner that aligns as analogical as possible with recent observations. Given that stochastic effects are more pronounced at higher redshifts, $\sigma_{\rm MAR}$ is assumed to be within $[-0.3,0.3]$ ($[-0.1,0.1]$) at $z\sim 9$ ($z<8$). The corresponding variation in absolute magnitude is $\sigma_{M_{\rm UV}}\sim 0.75\,(0.25)$, consistent with the variability obtained by Shen et al. (2023), Kravtsov & Belokurov (2024), and Wang et al. (2024).
@@ 2.5 Bayesian analysis
where $(x_{i},y_{i})$ are the observational data points, $f(x_{i})$ represents the predicted values, $d_{i}$ serves as a scale parameter, and $c_{i}$ is the asymmetry parameter. The priors for all primary parameters are detailed in Table 1. For an optimal balance between accuracy and efficiency, we implemented the nested sampling method, utilizing Nessai for Bayesian parameter estimation. We configured the analysis with 1000 live points [^4] Preliminary tests with 500 live points yielded identical results. Given the recommendation to set the number of live points higher than the dimensionality of the parameter space (Ashton et al., 2022), our configuration ensures adequate convergence. and an evidence tolerance of 0.5 to terminate the sampling process.
@@
Table 1: Prior distributions of the parameters for SFE, stochasticity and cosmology
@@
Figure 2: The left panel shows the observations of UV LFs (extinction corrected) and the best-fit results in the redshift range of $4\leq z\leq 10$. The right panel shows the corresponding distributions of SFE. The colored regions are at $68\%$ credible level. The light orange region represents the range of maximum value (20%-30%) of SFE (Wechsler & Tinker, 2018). Due to the redshift-dependent nature of SFE, which connects all observations at each redshift, the estimated UV LFs are tightly constrained.
[…]
@@ 3.1 The Evolution of SFE and UV LFs
As shown in Figure 2, we present the best-fit and posterior results of UV LFs and SFE, alongside the observations. The corresponding posterior distributions are shown in Figure 9555It is worth noticing that we have tested the assumption that the SFE do not evolve with redshift ($par_{s}=0$). The (logarithmic) Bayes factor of the redshift-evolved SFE scenario compared to the non-evolving scenario is $\ln{\mathcal{B}}=396.2$, which is strongly in favor of the redshift-evolution of the SFE.. For a better view, we present the UV LFs rather than the numbers of galaxies that are used practically. It is evident that the UV LF model fits well with the observations except for the data points at the bright end ($M_{\rm UV}<-23\rm mag$), which may be attributed to the assumption of dust extinction or the extra contribution of active galactic nuclear (AGN) in addition to the galaxy UV LFs. At the redshift range of $4\geq z\geq 7$, Harikane et al. (2022) confirmed that the AGN LFs dominate at $M_{\rm UV}<-24\,\rm mag$ and the galaxy LFs dominate at $M_{\rm UV}>-22\,\rm mag$. They also found that the bright-end excess results in galaxy UV LFs is better fitted with the double-power-law function rather than the classical Schechter function. One possible explanation for the discrepancy is the empirical $\beta_{\rm UV}-M_{\rm UV}$ relation, which is not well-constrained at very bright magnitudes. Additionally, the true attenuation curve for high-redshift galaxies remains poorly understood (Ma et al., 2019; Fudamoto et al., 2020; Liang et al., 2021). As a result, our model may have overestimated the dust obscuration, leading to larger uncertainties at the bright end. Furthermore, Harikane et al. (2022) also suggested that even after subtracting the number density of AGNs based on spectroscopic galaxy fractions, hidden AGNs could still be present within the galaxy LFs, further contributing to the observed bright-end excess. Since our model relies on the shape of SFE, it is more analogous to the Schechter function and shows slight inconsistency with the observations at the bright end. Besides, a fraction of data shows better fitting extent with the empirical model at $z=4$, which arises from their smaller measuring uncertainties and the more harmonious tendencies towards higher-redshift results.
The corresponding distributions of the evolved SFE are consistent with other researches (such as Fig. 2 in Wechsler & Tinker (2018)). The SFE peaks at $M_{1}\sim 10^{11.79}M_{\odot}$ for $z=4$ and $10^{12.84}\,\rm M_{\odot}$ for $z=10$ (note that the latter is only obtained for extrapolation since the extremely bright end of the UV LFs has not been detected at $z\geq 8$). At low halo masses, the profile of SFE slope is determined by $\beta$, which remains almost constant with $\beta\sim 0.50$ across wide redshift range. Similar conclusions are found in Feldmann et al. (2024), also. Conversely, at high masses, $\gamma$ dominates the profile which varies from $\sim 0.43$ to $\sim 1.34$, resulting in a notable steepening of the SFE. However, at $z\geq 8$, due to the absence of UV LFs at the bright end, the SFE profile at $M_{1}\geq 10^{12}\,\rm M_{\odot}$ can not be well constrained. Consequently, it remains inconclusive whether there is a reliable evolution of the SFE at high masses. As an extra supplementary, Figure 3 illustrates the stochastic effect referred in subsection 2.4. The posterior distributions of the DM halo accretion rate $\dot{M}_{\rm h}$ are consistent with the prior assumption, which is in line with the expectation. To further validate the reliability of our estimated results, we present a simplified comparison between our findings and those from GALLUMI. Detailed discussions are provided in Appendix C.
@@
where $M_{\rm trunc}$ denotes the truncation magnitude of the UV LFs. In general, $M_{\rm trunc}=-17$ corresponds to the SFR of $0.3\,\rm M_{\odot}\,yr^{-1}$ (Finkelstein et al., 2015; Oesch et al., 2018; Bouwens et al., 2020; Harikane et al., 2022). Additionally, we consider another truncated limit with $M_{\rm trunc}=-15$ following McLeod et al. (2016) and Ishigaki et al. (2018). As shown in Figure 4, the UV luminosity densities derived from the fitted UV LFs are in accordance with other observations. The empirical equation Eq.(15) in Madau & Dickinson (2014) fits well with the observations at low redshifts (i.e., $z\leq 6.5$) but tends to overestimate the UV luminosity density and the SFR density at higher redshifts.